RHIC Startup 09

ver  February 10, 2009 pws

Choose from:

GO to Gantt chart of overall start up

 

Critical Items:

 

RHIC cool-down completed

Completion of power supply checkout

Completion of DX training

Set up Yellow injection

Bring STAR magnets on for ramp development

 

SCHEDULE FOR LINAC:

 

·        Continue to run for BLIP and AGS

·        Dummy load for BM2 ps testing?

 

SCHEDULE FOR BOOSTER:

 

·        Continue to run for AGS

 

SCHEUDLE FOR AGS:

 

·        Continue to run for RHIC

 

RHIC SCHEDULE:

 

·        2/9    

o       Yellow cool-down complete

o       Begin Yellow ps checkout

o       Continue injection setup in Blue

·        2/11  

o       Complete Yellow ring power supply checkout

o       Begin work with beam in the Yellow Ring

 

 

Back to top

 

LINAC

 

·        October and November:

o       Complete outstanding shutdown work

 

Booster:

·        Early August

o       Install Dipoles and beam line components for EBIS in Booster tunnel In Progress

·        Late August

o       Bake out of TTB and LTB/EBIS transport line In Progress

o       Install HEBT cap In Progress

·        August 30 the Booster will be on Controlled Access in the evenings as needed for ACS re-certification

·        September 10-12:

o       Begin systems turn on for Booster

o       Remove Shutdown LOTO for Booster

o       Begin RF restoration

o       TTB power supply checkout

o       Sweep secure Tandem/AGS alcove

o       Begin chipmunk calibration for Booster

o       Tandem/TTB ACS re-certification

·        FRIDAY the 12th 1400-1700

o       Booster Hi-pot - NO RING ACCESS

·        September 17th 0700-1900

o       Booster MMPS checkout, ACS testing and Heat run - NO RING ACCESS

·        General work September 15-19

o       Begin active restoration of Booster systems

o       Daytime presence in MCR

o       Inspect Booster shielding and Berm

o       Complete chipmunk Booster calibration

o       Sweep/secure Booster, close plug door if practical

o       Begin pre-check of Booster, BtA and NSRL power supplies

o       Complete ACS re-certification

o       Begin RSC checklists for TTB, Booster BtA and NSRL

·        September 19th 0800-1400

o       Booster Dry run - NO RING ACCESS

·        September 22 AM ACCESS

o       Remove Vacuum components

o       Repair water leak at C2

o       Back flush D6

o       Investigate problems from Friday’s running

o       Load all functions for Fe 1GeV

o       Repair NSRL Berm gate

o       MCR Checkout

o       Begin Power supply and magnet checkout

·        September 23 VERY LIMTED RING ACCESS

o       Execute MCR checkout of TTB and Booster devices

o       Booster final inspection

o       Sweep/secure Booster Berm

o       Complete any outstanding Booster Jobs

o       Complete RSC checklists for TTB, Booster and NSRL

o       Transport Fe beam to TTB Beam stops

·        September 24 0800

o       Begin work with beam in the Booster – NO RING ACCESS TO BOOSTER, TTB CROSSOVER, TANDEM ALCOVE

·        November 10

o       1500Hrs- LOTO MMPS and Back-legs (930U-G and 914-1-2A) for vacuum work

·        November 21st NSRL ENDS

·        November 24th

o       0700-1000 Injection Bump testing- No ACCESS

o       Booster opened to RA following Surveys

·        Dec 8th

o       Begin work on C6 injection foil replacement

·        December 1-10

o       Begin LINAC startup for BLIP (and PP)

·        December 15

o       Sweep secure LINAC – LIMITED ACCESS

o       LINAC on for BLIP (and PP)

 

AGS

 

·        Tuesday Sept 30 0800-1600 AGS ring secured for power supply tests

·        October6-10

o       Restore vacuum to AGS

o       Check out L20, A5, F6, F3 and BXT bumps

o       Continue / complete ACS re-certification

o       Check out L20 angle and position

o       Close AGS Sleeve boxes

o       Remove Shutdown LOTO for instrumentation

o       Test / open AGS vacuum valves

o       Replace BtA flag at “150”

o       Reinstall and inspect shielding , D line AGS and Switchyard

·        October 13 0800-1500

o       Complete BtA flag replacement

o       Begin re-calibration BtA quads on AGS side

·        October 15th Begin BtA study

 

·        11/4

o       Complete injection studies

o       Apply RA LOTO to AGS

o       Use Alternate LOTO for AMMPS

 

·        11/5

o       Ring remains closed until Hi-Pot of AGS cold snake complete (~1100hrs)

o       Access ring after Hi-Pot for HP survey (CA, HP Access only)

o       ~1300 AGS to Restricted Access

o       Remove jumper on J10 water flow interlock

o       Remove alternate supply for AGS MMPS

·        11/6

o       Apply standard LOTO to AGS MMPS

o       Remove alternate LOTO when standard in place

o       Begin AGS A20 snake cool down to 80K

o       Begin AGS A20 snake cool down to 4k

·        11/17

o       Morning Sweep and secure AGS

o       Begin Siemens transformer testing and commissioning – NO ACCESS

·        12/5

o       Apply AGS RA LOTO

o       AGS to Restricted Access

 

RHIC

 

o       Start scrub of RHIC ring (21 days)

o       RHIC ring insulating vacuum established

o       Start 80K cool down of RHIC ring (4 weeks)

·        December 8th*

o       Start cryo scrub (plant)

·        Mid-Late December*

o       Start cryo scrub (ring)

o       45K cook down of cryo plant and RHIC ring

o       Hi-pot RHIC magnets at 45K

 

Back to top

 

Outline of startup 09:

 

Figure 1

Figure 2

Figure 3

 

Back to top